OpenStudy (anonymous):

Solve the equation x2 + 5x = 0.

OpenStudy (zzr0ck3r):

factor an x out and what do you have?

OpenStudy (anonymous):

no its x^2 it didnt copy the squared

OpenStudy (zzr0ck3r):

?

OpenStudy (zzr0ck3r):

yes I know

OpenStudy (zzr0ck3r):

\(x^2+5x=0\iff x(x+5)=0\)

OpenStudy (zzr0ck3r):

what does this tell us about x?

OpenStudy (anonymous):

wait now i just use the zero product property?

OpenStudy (zzr0ck3r):

yes

OpenStudy (zzr0ck3r):

what do you get?

OpenStudy (anonymous):

so x= 1,-5?

OpenStudy (zzr0ck3r):

\(x(x+5)=0 \iff x=0 \text{ and } x+5=0 \iff x=0 \text{ and } x=-5\)

OpenStudy (anonymous):

thank you

OpenStudy (zzr0ck3r):

np